# Causal structure and topology change in (2+1)-dimensional simplicial gravity

### v1

@@ 4. Indistinguishable lightcones:
(a) Since the uppermost triangle $\Delta(v_{2}v_{3}v_{4})$ is Minkowskian, the timelike region lies in both the past- and future-oriented lightcones.
@@ 5.1 Examples: Deficit angles and vertex causal structure
$S_{\rm Regge}({\cal T}_{1\text{-}4})=-\sum_{i=1}^{4}\sqrt{|s_{i}|}\,\delta_{0i}(s)-\sum_{i<j}\sqrt{|s_{ij}|}\,\delta_{ij}(s)\,,\quad$ (11)
@@
To investigate this, we examine three illustrative examples in which we fix the boundary data ${\cal S}_{\partial}=(s_{12},s_{13},s_{14},s_{23},s_{24},s_{34})$ and vary the bulk squared edge lengths $s_{i}=a$, for $i=1,\dots,4$ uniformly. For each example, we analyze the behaviour of the bulk deficit angles and the Regge action as functions of $a$, while simultaneously tracking the causal structure for the bulk vertex $v_{0}$ via $(N_{\rm sl},N_{\rm tl})$. Visualizations of the lightcone intersection patterns on the triangulated boundary sphere $\partial{\rm St}(v_{0})$ for each example are provided in Appendix B.
@@
We begin with a case in which the boundary data corresponds to a Euclidean geometry. Specifically, we choose
${\cal S}_{\partial 1}=(1,1,2,1,2,2)$
which satisfy the Euclidean simplex inequalities for the boundary triangulation $\partial\mathrm{St}(v_{0})$. For this choice of boundary data, the four tetrahedra of the complex ${\cal T}_{1\text{-}4}$ satisfy the full set of Lorentzian simplex inequalities provided the common bulk squared edge length satisfies $a<1/3$. The infinite range for the bulk lengths is reminiscence of spike configurations studied in [55].
Figure 11 plots the bulk deficit angle $\delta_{01}$ and the Regge action $S_{\rm Regge}$ as functions of $a$. Two qualitative transitions occur at $a=0$ and $a=1/4$, where the imaginary part of the deficit angles $\delta_{0i}$ and the Regge action exhibit singularities. The corresponding change in vertex causal structure for the different intervals are summarized below and visualized in Figure 14 of Appendix B.
@@
Figure 11: Bulk deficit angle $\delta_{01}$ and Regge action $S_{\rm Regge}$ versus the squared bulk length $a$ for the Euclidean boundary data ${\cal S}_{\partial 1}$. The green step function indicates the range in which all four tetrahedra satisfy the Lorentzian simplex inequalities. Dashed lines represent imigainary parts. The vertex causal structure transitions at $a=0$ and $a=\tfrac{1}{4}$ where the deficit angles and the Regge action become singular.
@@
This example highlights the independence of edge-causality and vertex-causality. In Regions I and III, there exist edges with a causally regular structure but the bulk vertex structure is irregular.
Example 2.
As a second illustration, we consider the boundary geometry that is intrinsically Lorentzian with
${\cal S}_{\partial 2}=(3,3,2,-3,2,2)$
satisfying Lorentzian simplex inequalities for the boundary tetrahedron $\Delta(v_{1}v_{2}v_{3}v_{4})$. For this configuration, the triangle inequalities for the four bulk tetrahedra are satisfied in the finite interval 3/$53/5<a<4/$ 5.
Refer to captionRefer to caption
Figure 12: Bulk deficit angle $\delta_{01}$ and Regge action $S_{\rm Regge}$ as functions of the squared bulk length $a$ for Lorentzian boundary data ${\cal S}_{\partial 2}$. Both quantities remain real in the admissible range 3/$53/5<a<4/$ 5. The right panel provides a zoomed-in view of the Regge action to show its smooth behaviour.
@@
[…]
which yield a Gram matrix with signature $(-1,-1,+1)$. This violates both the Euclidean and Lorentzian simplex inequalities. Nevertheless, all four bulk tetrahedra in the complex ${\cal T}_{1\text{-}4}$ satisfy the Lorentzian simplex inequalities for squared bulk edge lengths $a>1$. This provides an example of a valid Lorentzian simplicial 3-complex with unphysical boundary geometry–i.e., not itself a triangulation of a valid tetrahedron.
@@
  • Singularities and Topology change: We observe that changes in vertex–causality (determined by transitions in the pair $(N_{\rm sl},N_{\rm tl})$) often coincide with singularities in the bulk deficit angles. These singularities signal changes in the lightcone structure and are typically associated with boundary faces becoming null. Such transitions may represent discrete analogues of Morse-type critical points where topology change can occur in the triangulated spacetime.
@@
  • Mismatch between causality notions: The examples demonstrate that vertex–causality and hinge–causality encode different aspects of the local light-cone structure. The observed divergences in curvature and action at transitions in vertex–causality can be viewed as the appearance or resolution of analogues of point-like singularities in the discrete spacetime. A causally regular vertex does not guarantee regularity at the adjacent edges/hinges, and vice versa. For instance, Example 1 (for $a<0$) provides a configuration where the bulk vertex has an irregular structure but the incident edges remain causally regular [^12] Note that it is possible to find configurations violating causal regularity at the vertex while all incident edges remain causally regular, including configurations with spacelike bulk edges. Such cases can arise in larger triangulations. See [64] for details.. See [16, 12] for a similar configuration in a four dimensional triangulation. In contrast, Example 3 in the interval $41<a<5/$ 4 shows a configuration where the vertex is causally regular but there are causally irregular edges incident on it.
@@ 6 Discussion
Importantly, our classification enables a refined understanding of causally irregular configurations within simplicial quantum gravity. In analogy with continuum geometry, causally irregular edges, i.e., hinges with nontrivial lightcone crossings, can be interpreted as discrete analogues of conical singularities, where curvature and causal structure concentrate along codimension-2 defects. Likewise, causally irregular vertices, where the lightcone intersection pattern becomes topologically nontrivial, correspond to point-like causal singularities, representing localized breakdowns of Lorentzian structure in the simplicial manifold. These distinctions provide useful tools for identifying potential sites of signature change, topology change, or non-classical behavior in path integral configurations.
@@
The methods developed in this analysis have broader significance beyond the setting of 2+1 dimensions. In fact, this study automatically captures the causal structure associated with codimension-3 sub-simplices in higher-dimensional triangulations. For example, in 3+1 dimensions, the sub-simplices of codimension-3 are edges. Projecting out such edges yields a simplicial neighbourhood that is topologically homeomorphic to a three-ball, just like a vertex neighbourhood in 2+1 dimensions. Thus, the procedures developed here naturally extend to characterizing hinge–causality in (3+1)–dimensional triangulations.
Understanding these singular structures, both conical and point-like, could be essential for identifying and regulating the contributions of non-classical geometries in quantum gravity path integrals. For example, one might use causal regularity criteria to define improved measures or selection rules that suppress configurations with pathological singularities, or conversely, to probe regions of the configuration space relevant to non-perturbative phenomena such as wormholes, causal fluctuations, or signature transitions.
The procedures employed in this work also extend naturally to the study of vertex–causality in higher dimensions. The construction of local simplicial neighoburhoods, the embedding of a simplex in $\mathbb{R}^{1,d-1}$ from its edge lengths, and classification of the lightcone intersection patterns are all geometrically well-defined in arbitrary dimensions. In higher dimensions, the intersection of the lightcones with the boundary star of a vertex yields a collection of null hypersurfaces embedded in the boundary ($d-$1)–simplices. We leave a detailed study of vertex–causality in $d\geq 4$ for future work.
@@ Acknowledgements
The authors thank Sebastian Steinhaus and the Emmy Noether junior research group at FSU Jena for many fruitful discussions. The authors also thank Martin Ammon and Bianca Dittrich for helpful discussions. SKA gratefully acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) project number 422809950.
@@ Appendix A Proof of realization of simplex in Minkowski spacetime
Let $G$ be a real symmetric $d\times d$ matrix with signature (,+,+)(-,$\dots$,+)( - , + … , + ). By the spectral theorem, $G$ is diagonalizable, i.e., there exist an orthogonal matrix $Q\in O(d)$ and a diagonal matrix $\Lambda$ such that $G=Q\Lambda Q^{\top}$. W.l.o.g., we assume the negative eigenvalue to be $\Lambda_{11}$, else reorder the basis. Now, define the realization matrix
@@
using $\lvert\Lambda\rvert_{ij}\coloneqq\lvert\Lambda_{ij}\rvert$ to ensure the existence of a real square root. By construction, $\lvert\Lambda\rvert\eta=\Lambda$ holds, where $\eta={\rm diag}(-1,1,\dots,1)$ is the Minkowski metric. To prove that $L$ is indeed the realization matrix which contains the edge vectors $\{e_{i}\}$ of a simplex, we show:
@@
Appendix B Snapshots of vertex causality for the examples
This appendix supplements Section 5 by displaying, for each example, the lightcone intersection patterns on the boundary 2-sphere $\partial\mathrm{St}(v_{0})$. Each pattern is shown as a planar “unfolding” of the four boundary triangles; folding the diagram back to a tetrahedron immediately yields the invariants $(N_{\rm sl},N_{\rm tl})$.
Example 1:
In Example 1, the vertex causal structure transitions at $a=0$ and $a=\tfrac{1}{4}$. Figure 14 displays lightcone intersection patterns on the unfolded 2-sphere $\partial\mathrm{St}(v_{0})$ in each interval.
  • Region I $(a<0)$: All four tetrahedra within the triangulation ${\cal T}_{\text{1-4}}$ are of type A1, resulting in a causally irregular vertex with $(N_{\rm sl},N_{\rm tl})=(0,1)$.
  • Region II $(0<a<1/4)$: All the tetrahedra in ${\cal T}_{\text{1-4}}$ are of type D5, resulting in a irregular vertex causal structure with $(N_{\rm sl},N_{\rm tl})=(4,1)$. Each vertex $v_{i}$ is contained in one of the four disjoint spacelike regions.
  • Region III $(1/4<a<1/3)$: Here, $v_{0}$ also has an irregular causal structure with $(N_{\rm sl},N_{\rm tl})=(2,2)$. The spacelike region containing the vertices $v_{1},v_{2}$ and $v_{3}$ is disconnected from the spacelike region containing the vertex $v_{4}$.
[…]
Figure 14: Lightcone intersection patterns on the boundary 2-sphere for Example 1. The shaded domains are spacelike, white domains are timelike, and red curves are null intersections.
Example 2:
Here, the simplex inequalities for the entire triangulation are satisfied in a finite range 3/$53/5<a<4/$ 5. The bulk vertex $v_{0}$ has a regular structure with $(N_{\rm sl},N_{\rm tl})=(1,2)$, as depicted in Figure 15.
$v_{4}$$v_{4}$$v_{4}$$v_{2}$$v_{1}$$v_{3}$
Figure 15: Snapshot of causal structure for Example 2. The lightcone intersection patterns on the boundary sphere shows a regular causal structure.
Example 3:
In this example, the causal structure of the vertex $v_{0}$ transitions between the two intervals allowed by the simplex inequalities. Figure 16 shows the lightcone structure on the boundary 2-sphere which transitions from a causal regular configuration in the interval $41<a<5/$ 4 to an irregular configuration (acausal) in the interval $a>5/4$.
$v_{4}$$v_{4}$$v_{4}$$v_{2}$$v_{1}$$v_{3}$$v_{4}$$v_{4}$$v_{4}$$v_{2}$$v_{1}$$v_{3}$
Figure 16: Lightcone intersection patterns for Example 3. The boundary sphere transitions from a regular structure to an acausal one.

### v2

@@ 4. Indistinguishable lightcones:
(a) Since the uppermost triangle $\Delta(v_{2}v_{3}v_{4})$ is Lorentzian, the timelike region lies in both the past- and future-oriented lightcones.
@@ 5.1 Examples: Deficit angles and vertex causal structure
$S_{\rm Regge}({\cal T}_{1\text{-}4})=-\sum_{i=1}^{4}\sqrt{|s_{i}|}\,\delta_{0i}(s)-\sum_{i<j}\sqrt{|s_{ij}|}\,\delta_{ij}(s)\,$$\quad$ (14)
@@
To investigate this, we examine three illustrative examples in which we fix the boundary data ${\cal S}_{\partial}=(s_{12},s_{13},s_{14},s_{23},s_{24},s_{34})$ and vary the bulk squared edge lengths $s_{i}=a$, for $i=1,\dots,4$ uniformly. For each example, we analyze the behaviour of the bulk deficit angles and the Regge action as functions of $a$, while simultaneously tracking the causal structure for the bulk vertex $v_{0}$ via $(N_{\rm sl},N_{\rm tl})$. We provide visualizations of the lightcone intersection patterns on the triangulated boundary 2-sphere $\partial{\rm St}(v_{0})$ for each example. Each pattern is represented as a planar unfolding of the four boundary triangles, from which the invariants $(N_{\rm sl},N_{\rm tl})$ can be inferred by conceptually refolding the triangles to reconstruct the triangulated boundary 2-sphere or tetrahedron.
Before turning to explicit examples, it is important to clarify the nature of the boundary complex of the triangulation ${\cal T}_{1\text{-}4}$. Although this boundary complex is topologically equivalent to a tetrahedral surface, its geometry does not necessarily coincide with that of a single flat tetrahedron [^12] In cases where the boundary tetrahedron is geometrically flat and the bulk vertex $v_{0}$ lies inside it, the full triangulation ${\cal T}_{1\text{-}4}$ can be embedded in $\mathbb{R}^{1,2}$.. The individual boundary triangles are glued along shared edges of adjacent simplices, but the resulting boundary geometry may not embed globally as a single tetrahedron in $\mathbb{R}^{1,2}$. The examples considered below represent consistent simplicial configurations with non-degenerate local geometry, but with distinct geometric realizations of the boundary data. Our focus will be on analyzing the bulk vertex–causality in each case.
@@
We begin with an example in which the boundary geometry is Lorentzian. The boundary data is chosen as
${\cal S}_{\partial 1}=(3,3,2,-3,2,2),$
and satisfies the Lorentzian simplex inequalities for the boundary tetrahedron $\Delta(v_{1}v_{2}v_{3}v_{4})$, i.e., the corresponding Gram matrix has signature (++)(-++). For this configuration, the triangle inequalities for all the four bulk tetrahedra are satisfied for the bulk edge length in the finite interval $3/5<a<4/5$.
Refer to captionRefer to caption
Figure 11: Bulk deficit angle $\delta_{01}$ and Regge action $S_{\rm Regge}$ as functions of the squared bulk length $a$ for Lorentzian boundary data ${\cal S}_{\partial 2}$. Both quantities remain real in the admissible range $3/5<a<4/5$. The right panel provides a zoomed-in view of the Regge action to show its smooth behaviour.
$v_{4}$$v_{4}$$v_{4}$$v_{2}$$v_{1}$$v_{3}$
Figure 12: Snapshot of the vertex causal structure for Example 1 with boundary data ${\cal S}_{\partial 1}$ within the region $3/5<a<4/5$. The lightcone intersection patterns on the boundary sphere shows a regular causal structure.
Example 2.
We next consider an example where the boundary data corresponds to a Euclidean tetrahedron geometry. Specifically, we choose
${\cal S}_{\partial 2}=(1,1,2,1,2,2)$
which satisfy the Euclidean simplex inequalities for the boundary triangulation $\partial\mathrm{St}(v_{0})$, i.e., the corresponding Gram matrix has signature $(+,+,+)$. This indicates that the boundary tetrahedron is embeddable in Euclidean space, and hence the complex ${\cal T}_{1\text{-}4}$ cannot be embedded in $\mathbb{R}^{1,2}$. For this choice of boundary data, each of the four tetrahedra of the complex ${\cal T}_{1\text{-}4}$ satisfy the Lorentzian simplex inequalities provided the common bulk squared edge length satisfies $a<1/3$. This infinite range for the bulk lengths is reminiscence of spike configurations studied in [55].
Figure 13 plots the bulk deficit angle $\delta_{01}$ and the Regge action $S_{\rm Regge}$ as functions of $a$. Two qualitative transitions occur at $a=0$ and $a=1/4$, where the imaginary parts of the dihedral angles $\delta_{0i}$ are discontinuous and the Regge action becomes non-differentiable. The corresponding change in vertex causal structure for the different intervals are summarized below and visualized in Figure 14.
@@
Figure 13: Bulk deficit angle $\delta_{01}$ and Regge action $S_{\rm Regge}$ versus the squared bulk length $a$ for the Euclidean boundary data ${\cal S}_{\partial 1}$. The green step function indicates the range in which all four tetrahedra satisfy the Lorentzian simplex inequalities. Dashed lines represent imaginary parts. The vertex causal structure transitions at $a=0$ and $a=\tfrac{1}{4}$.
@@
$v_{4}$$v_{4}$$v_{4}$$v_{2}$$v_{1}$$v_{3}$Region $v4v_{4}v4v_{4}v4v_{4}v2v_{2}v1v_{1}v3v_{3}Region III$
Figure 14: Snapshots of the causal structure of the bulk vertex in Example 2 for the causally distinct intervals. The grey regions are spacelike, white regions are timelike, and red curves are lightcone intersections with respect to the bulk vertex. Region I lies in the interval $a<0$, Region II in $0<a<\tfrac{1}{4}$ and Region III in $\tfrac{1}{4}<a<\tfrac{1}{3}$.
This example highlights the independence of edge-causality and vertex-causality. In Regions I and III, there exist bulk edges with causally regular structures but the bulk vertex structure is irregular.
@@
${\cal S}_{\partial 3}=(5,\,5,\,-5,\,-5,\,-1,\,-1),$
which yield a Gram matrix with signature (+)(-\,-\,+). This violates both the Euclidean and Lorentzian simplex inequalities [^13] The boundary tetrahedron with this data is only embeddable in $\mathbb{R}^{2,1}$ with two time coordinates. . Nevertheless, all four bulk tetrahedra in the complex ${\cal T}_{1\text{-}4}$ satisfy the Lorentzian simplex inequalities for squared bulk edge lengths $a>1$.
@@
$v_{4}$$v_{4}$$v_{4}$$v_{2}$$v_{1}$$v_{3}$$v_{4}$$v_{4}$$v_{4}$$v_{2}$$v_{1}$$v_{3}$
Figure 16: Snapshots of the lightcone intersection patterns for Example 3 on the boundary 2-sphere. The causal structure of the bulk vertex $v_{0}$ transitions from a regular to an acausal configuration.
@@
  • Discontinuities and Topology change: We observe that changes in vertex–causality (determined by transitions in the pair $(N_{\rm sl},N_{\rm tl})$) often coincide with discontinuities or kinks in the bulk deficit angles. These discontinuities signal changes in the lightcone structure and are typically associated with boundary faces becoming null. Such transitions may represent discrete analogues of Morse-type critical points where topology change can occur in the triangulated spacetime.
@@
  • Mismatch between causality notions: The examples demonstrate that vertex–causality and hinge–causality encode different aspects of the local light-cone structure. The observed divergences in curvature and action at transitions in vertex–causality can be viewed as the appearance or resolution of analogues of point-like singularities in the discrete spacetime. A causally regular vertex does not guarantee regularity at the adjacent edges/hinges, and vice versa. For instance, Example 1 (for $a<0$) provides a configuration where the bulk vertex has an irregular structure but the incident edges remain causally regular [^14] Note that it is possible to find configurations violating causal regularity at the vertex while all incident edges remain causally regular, including configurations with spacelike bulk edges. Such cases can arise in larger triangulations. See [64] for details.. See [16, 12] for a similar configuration in a four dimensional triangulation. In contrast, Example 3 in the interval $1<a<5/4$ shows a configuration where the vertex is causally regular but there are causally irregular edges incident on it.
@@ 6 Discussion
Importantly, our causal classification refines the understanding of causal irregularities in simplicial quantum gravity. In direct analogy with the continuum, where curvature and causal structure can become singular along lower-dimensional defects, hinge–causality and vertex–causality capture their discrete counterparts. Causally irregular hinges (edges with non-canonical causal structure in their projected local neighbourhood) act as discrete analogues of conical singularities, representing curvature localized along codimension–2 defects. Similarly, causally irregular vertices, where the lightcone intersections on the boundary sphere acquire a non-canonical configuration, correspond to point-like causal singularities that signal a breakdown of the local lightcone structure in the simplicial manifold. Together, these distinctions provide a geometric framework for identifying discrete analogues of topology change, and other forms of non-classical phenomena that may contribute non-trivially to the gravitational path integral.
@@
The methods developed in this analysis have broader significance beyond the setting of 2+1 dimensions. In fact, this study automatically captures the causal structure associated with codimension–3 subsimplices in higher-dimensional triangulations. For example, in 3+1 dimensions, the subsimplices of codimension–3 are edges. Projecting out such edges yields a simplicial neighbourhood that is topologically homeomorphic to a three-ball, just like a vertex neighbourhood in 2+1 dimensions. Thus, the procedures developed here naturally extend to characterizing hinge–causality in (3+1)–dimensional triangulations.
Understanding these singular structures, both conical and point-like defects, could be essential for identifying and regulating the contributions of non-classical geometries in quantum gravity path integrals. For example, one might use causal regularity criteria to define improved measures or selection rules that suppress configurations with pathological singularities, or conversely, to probe regions of the configuration space relevant to non-perturbative phenomena such as wormholes, causal fluctuations, or signature transitions. Notably, in the context of topology change in two dimensions, it has been shown that introducing complexified metrics in the continuum can regularize curvature and causal singularities [14, 53]. It would be interesting to extend such analysis to higher dimensions and explore whether complex metrics can similarly account for the point-like causal irregularities identified in this work.
The procedures employed in this work also extend naturally to the study of vertex–causality in higher dimensions. The construction of local simplicial neighbourhoods, the embedding of a simplex in $\mathbb{R}^{1,d-1}$ from its edge lengths, and classification of the lightcone intersection patterns are all geometrically well-defined in arbitrary dimensions. In higher dimensions, the intersection of the lightcones with the boundary star of a vertex yields a collection of null hypersurfaces embedded in the boundary (dd-1)–simplices. We leave a detailed study of vertex–causality in $d\geq 4$ for future work.
@@ Acknowledgements
The authors thank Sebastian Steinhaus and the Emmy Noether junior research group at FSU–Jena for many fruitful discussions. We are also grateful to Martin Ammon and Bianca Dittrich for helpful discussions, and to the anonymous referee for constructive suggestions that improved clarity of the manuscript. SKA gratefully acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) project number 422809950.
@@ Appendix A Proof of realization of simplex in Minkowski spacetime
Let $G$ be a real symmetric $d\times d$ matrix with signature $(-,+\dots,+)$. By the spectral theorem, $G$ is diagonalizable, i.e., there exist an orthogonal matrix $Q\in O(d)$ and a diagonal matrix $\Lambda$ such that $G=Q\Lambda Q^{\top}$. W.l.o.g., we assume the negative eigenvalue to be $\Lambda_{11}$, else reorder the basis. Now, define the realization matrix
@@
using $\lvert\Lambda\rvert_{ij}\coloneqq\lvert\Lambda_{ij}\rvert$ to ensure the existence of a real square root. By construction, $\lvert\Lambda\rvert\eta=\Lambda$ holds, where $\eta={\rm diag}(-1,1,\dots,1)$ is the Minkowski metric. To prove that $L$ is the realization matrix which contains the edge vectors $\{e_{i}\}$ of a simplex, we show: